# E4S: Fine-grained Face Swapping via Editing With Regional GAN Inversion

### v2

@@ 3.1.1 Cropping and Reenactment
3.1.2 Swapping and Generation
After the cropping and reenactment process, we are ready to fulfil the face swapping process. We first feed the driven pair ($I_{d},M_{\rm{d}}$) and the target pair ($I_{\rm{t}},M_{\rm{t}}$) into a mask-guided multi-scale encoder $F_{\phi}$ respectively, obtaining the style codes to represent the texture of each facial region. This step can be described as:
$S_{t}=F_{\phi}(I_{\rm{t}},M_{\rm{t}}),\quad S_{d}=F_{\phi}(I_{d},M_{\rm{d}}),$ (3)
where $S_{t}$ and $S_{d}$ denote the extracted texture codes of the target and driven face, respectively. The detailed modules of the encoder $F_{\phi}$ are introduced in Sec. 3.2. Then, we exchange the texture codes of several facial components of $S_{t}$ with those of $S_{d}$, resulting in the recomposed texture codes $S_{swap}$. The components that must be exchanged are: eyebrows, eyes, nose, mouth, lips, face, skin, neck, and ears. One may be confused about the exchange of skin, since this paper aims to maintain the skin tone and lighting of the target. The reason we use the skin component from the source is replacing the texture codes of skin from the source face with the target one can not bring satisfactory results. Thus, we turn to a skin recoloring network for post-processing, rather than keeping the skin color of the target directly in the swapping process.
Next, we feed the recomposed mask $M_{\text{swap}}$ and texture codes ${S_{\text{swap}}}$ into the StyleGAN generator $G_{\theta}$ with a mask-guided style injection module to synthesize the swapped face $\tilde{I}$, which can be described as:
$\tilde{I}=G_{\theta}(M_{\text{swap}},S_{\text{swap}}).$ (4)
The details of the generator $G_{\theta}$ will be introduced in the Sec. 3.2. Note that our E4S does not need to train an extra sub-network to deal with the occlusion and can naturally achieve more accurate occlusion recovery than the existing methods FSGAN [15] and FaceShifter [9]. This is because the generator $G_{\theta}$ can fill out the occlusion pixels with the swapped texture features adaptively according to masks.
3.1.3 Post-Processing
Since the disentanglement of texture and shape solely cannot guarantee a lighting-natural swapped result, we additionally propose a face-recoloring network $B_{\psi}$ to conduct transferring of skin and lighting from the target face $I_{\rm{t}}$ to the swapped face $\tilde{I}$, which is elaborated in Sec. 3.3.1. Then, the swapped face $\tilde{I}$ and target image $T$ are blended together via Multi-Band Blending, producing the naive final swapped image $F_{final}$, which is detailed in Sec. 3.3.2. Finally, considering that the potential mismatch during the mask recombination process of $M_{\rm{d}}$ and $M_{\rm{t}}$ would cause difficulty for the shape keeping of source face, we propose an inpainting network $P_{\tau}$, which operates on pixel level and maintain the source face shape by the guide of mismatched region masks. we give the detailed description of the inpainting network $P_{\tau}$ in the Sec. 3.3.3.
3.2 Disentangling Shape and Texture
The key insight for the proposed E4S framework is disentangling the per-region texture and its corresponding shapes. To pursue a better disentanglement of shape and texture as well as the high-resolution and high-fidelity generation, we resort to the powerful generative model StyleGAN that can generate images with 1024$\times$1024 resolution naturally. Specifically, we aim to develop a GAN inversion method rather than training StyleGAN from scratch. That is, we leverage a pre-trained StyleGAN for the disentanglement, which also avoids training instability.
3.2.1 Mask-guided Style Extraction
Assuming there is an image $I$ and its segmentation mask $M$, we first leverage a multi-scale encoder $F_{\phi}$ to take the image $I$ as input and produce feature maps at different levels as:
@@ 3.2.1 Mask-guided Style Extraction
where $s_{j}$ indicates the style codes of the $j$-th facial region. Then, the generator $G_{\theta}$ takes as input the style codes and the mask $M$ are fed into the generator to generate the swapping result. Formally, we define $s\in\mathbb{R}^{C\times 18\times 512}$ as the proposed $\mathcal{W}^{r+}$ space.
Figure 4: Detailed comparison of the original StyleGAN and the proposed mask-guided StyleGAN. (a) The original StyleGAN contains consecutive convolution blocks. Each block contains a modulation, a demodulation, and a convolution layer. The $W$ and $b$ denote the learnable kernel weights for each block, and $s$ denotes the style code. The B stands for noise broadcast operation. An upsampling layer is used between every two blocks. (b) Our mask-guided StyleGAN regionally extends the convolution block. We sum up the intermediate feature maps of each region according to its segmentation mask which is downsized in advance.
3.2.2 Mask-guided Style Injection
The original StyleGAN contains a style-based generator taking 18 style codes with the dimension of 512 as input. The style codes are used to manipulate the feature maps of the 18 intermediate layers. As shown in Fig. 4 (a), the generator, consisting of a serial of style blocks, takes a constant feature map with the spatial size of 4$\times$4 as input. Each style block includes a modulation, a demodulation, 3$\times$3 convolution layer, and a noise layer B used to increase the diversity. Here, the learnable kernel weights and bias in each block are denoted as $W$ and $b$, respectively. $W$ would be scaled by its corresponding style code $s$ with the shape of 1$\times$512 before the convolution layer. An additional upsampling layer by the factor of two is adopted between every two style blocks, increasing the resolution of feature maps.
@@ 3.2.2 Mask-guided Style Injection
where $F_{l-1}$ and $F_{l}$ indicate the input and output feature maps of $l$-th layer, separately. $W^{\prime}_{jl}$ represents the scaled kernel weights for the $j$-th component in the $l$-th layer, and *** denotes the convolution operation. Similar to Eq. 6, the $\lfloor M\rfloor_{l}$ is the downsized mask corresponding to the $l$-th layer. Following the same modulation and demodulation as the original StyleGAN, we extend the style modulation regionally. In Eq. 9, $W_{l}$ denotes the original kernel weights for the $l$-th layer, and the $s_{jl}$ indicates the style code of $j$-th component for the $l$-th layer.
@@
The previous work SofGAN [46] is similar to our mask-guided StyleGAN since both of us regionally extend the style modulation. However, SofGAN applies an extra encoding module to map the mask $M$ to a 4$\times$4$\times$512 feature map. In contrast, we resort to the pretrained StyleGAN constant instead. Besides, SofGAN leverages spatially adaptive normalization [44] in the intermediate layers to strengthen the influence of the segmentation maps, while ours does not need since we find our region-aware convolution can already control the layout of each facial component well.
@@
Figure 5: Illustration of our re-coloring network $B_{\psi}$. In the training stage, we use two image augmentation to obtain the random re-colored $I^{\prime}_{\rm{AG}}$ and flipped $I^{\prime}_{\rm{A}}$. Then, we apply FPN [49] to acquire corresponding features $f_{\rm{AG}}$ and $f_{\rm{A}}$, where the feature of same semantic region in $f_{\rm{AG}}$ and $f_{\rm{A}}$ are used to calculate attention maps. Next, we multiply the maps with the downsampled $I^{\prime}_{A}$ to generate color guidance for each semantic region. Consequently, the composed color guidance $\Pi$, input images $I^{\prime}_{\rm{AG}}$ and $I^{\prime}_{\rm{A}}$ and their segmentation masks are concatenated through channel dimension and together fed to a U-Net to generate the re-colored result $\tilde{I}^{*}$. In the testing stage, given the swapped face $\tilde{I}$ to be recolored and the cropped target $I_{\rm{t}}$ providing color conditions, $B_{\psi}$ can transfer the color from $I_{\rm{t}}$ to $\tilde{I}$, resulting in the re-colored swapped result $\tilde{I}^{*}$.
3.3 Post-Processing for Natural Lighting and Face-Shape Consistency
Although the E4S framework can provide a unique face swapping insight (i.e, disentangling shape and texture), there are still some situations that it cannot handle well.
(i) When the lighting difference between source and target is huge, reconstructing the face of source in the ambient lighting of target would produce disharmonious results.
(ii) Directly using alpha-blending to blend the swapped face $\tilde{I}$ and target $T$ would lead to cliff edges, resulting in unnatural appearance in border area (e.g., hairline).
@@ 3.3 Post-Processing for Natural Lighting and Face-Shape Consistency
(iii) During the process of generating the recomposed mask $M_{\text{swap}}$, there would inevitably be some unmatched mask regions. For instance, for those positions belong to the inner face of the target, but not to that of the source, to maintain the shape of the source, these unmatched face regions should be injected with the style code of non-inner face category (hair, neck, etc.). However, different unmatched regions may correspond to different injected style codes. Furthermore, even if the mismatch regions are injected with the correct style code and yield a perfect recomposed mask $M_{\text{swap}}$, guided by this mask, the content generated in the mismatch regions tends to conflict with the original target image $T$. It brings difficulty for conducting blending on the reconstructed image $\tilde{I}$ and target $T$.
@@
To address the above issues, we additionally provide three post-processing techniques for more fine-grained and lighting-consistent face swapping, which consist of: transferring lighting condition and skin tones from the target to the swapped result with an additional face re-colorer $B_{\psi}$, Multi-Band Blending on the recolored image $\tilde{I}^{*}$ and non-cropped target $T$, and training an additional inpainting network $P_{\tau}$ to restore mismatch-mask regions.
3.3.1 Face Re-coloring
@@
Preserving the target lighting condition is a challenging task, especially when the illumination of the source is leaked into the result. Disentangling the illumination requires recovering material, geometry, and lighting of a scene, making this issue ill-posed. To deal with the re-lighting challenge, we reformulate the re-lighting problem as a re-coloring problem and propose to transfer the color from the target to the swapped face by our re-colorer $B_{\psi}$. Although the swapped face color should be determined by both the source skin tone and target environment lighting, we simplify the re-lighting problem by transferring the target color to the swapped face. In practice, we found this simplification leads to reasonable and admirable results, improving the result fidelity, as shown in Fig. 8.
@@
However, it is impossible to collect annotated paired data for the training of transferring the color of a reference face to the objective face. It may take great expense and huge labor to manually paint reasonable output. Thus, inspired by HeSer [23], we turn to perform self-supervised training, where we set the face to be colored and the reference face as the same identity during training. Specifically, for a given RGB face $I_{\rm{A}}$, we convert it into a grayscale image $I_{\rm{AG}}$. As seen in Fig. 5, we first conduct random color augmentation ${\rm{CA}}$ on the grayscale image $I_{\rm{AG}}$ and a random horizontal flip augmentation ${\rm{FA}}$ on the original image $I_{\rm{A}}$ respectively:
@@
$f_{\rm{AG}}={\rm{FPN}}({I_{\rm{AG}}^{\prime}})\in\mathbb{R}^{N\times C},$ (11)
$f_{\rm{A}}={\rm{FPN}}(I_{\rm{A}}^{\prime})\in\mathbb{R}^{N\times C},$
@@
Furthermore, since the high computational cost of calculating cosine similarity in 1024$\times$1024 resolution (e.g. calculating cosine similarity matrix between the skin parts of two 1024$\times$1024 faces requires 64GB memory in average), we train the U-Net in 256$\times$256 resolution. Although the memory is more efficient, the produced $256$ re-coloring result looks very blur when we directly resize it to 1024$\times$1024 resolution. Therefore, we use a face super-resolution method [50] to enhance the low-resolution result and paste the enhanced recolored image to $\tilde{I}$ based on a low-pass mask where the high-frequency pixels is removed by a Sobel filter, as shown in Fig. 8.
3.3.2 Multi-Band Blending
It is a natural choice to use Alpha Blending to blend the recolored swapped face $\tilde{I}^{*}$ and target image $T$ as the naive final swapped image $I_{\rm{\alpha}}$, which can be expressed as:
$I_{\rm{\alpha}}=\tilde{I}^{*}*M^{\rm{inner}}_{\rm{swap}}+T*(1-M^{\rm{inner}}_{\rm{swap}}),$ (15)
where $M^{\rm{inner}}_{\rm{swap}}$ indicates the inner face area of the swapped face $\tilde{I}^{*}$. However, such a result would have a stiff splice on the border of the inner face area (e.g., hairline). To deal with this, we use Multi-Band Blending [51, 52] to paste the swapped face $\tilde{I}^{*}$ back to the original target $T$. Multi-Band Blending conducts image smoothing in the low-frequency band and image sharpening in the high-frequency band. The Multi-Band Blending process can be formulated as:
$\tilde{I}^{*}_{\rm{paste}}={\rm{MBB}}(\tilde{I}^{*},T,M^{\rm{inner}}_{\rm{swap}})$ (16)
where ${\rm{MBB}}(\cdot)$ denotes the Multi-Band Blending algorithm. Concretely, a typical Multi-Band Blending algorithm consists of three steps: building image pyramids, blending computation, and reconstruction. Multi-Band Blending is able to seamlessly blend two images in a given mask region (i.e. $M^{\rm{inner}}_{\rm{swap}}$ in our case).
Figure 6: Illustration of inpainting network $P_{\tau}$, which adaptively inpaints the mismatch regions on the pixel levels. Given a driven face $I_{\rm{d}}$ and the pasted $\tilde{I}^{*}_{\rm{paste}}$, we first calculate the mismatch regions mask $M_{\rm{r}}$. Then the image $\tilde{I}^{*}_{\rm{paste}}$ along with the mismatch mask $M_{\rm{r}}$ are fed to an auto-encoder to inpaint the mismatch pixels. The generation process of the decoder is modulated by a vector $\alpha$ extracted from the area ratio $s$ of mismatch regions.
3.3.3 Inpainting For Mismatch-mask Region
There would inevitably be quite a few mismatched regions during facial mask exchange between the driven face $M_{\rm{d}}$ and target $M_{\rm{t}}$ (Sec. 3.1). Here, we divide facial mask into two categories: inner face region (eyebrows, eyes, nose, mouth, lips, face, and skin) and non-inner face region (neck, ears, hair, earrings, and background). Then, there are two types of face mismatched pixels in the recomposed mask $M_{\text{swap}}$: 1) the positions belong to the inner face of the source, but not to that of the target; 2) the positions belong to the inner face of the target, but not to that of the source. In the first case, to keep face shape of the source, we directly use face mask of the source to fill those mismatched positions. For the second case, we need to fill the position with a proper non-inner face category. In this paper, we define a mismatched region $M_{\rm{r}}$ as the second case since it cannot be solved directly:
@@ 3.3.3 Inpainting For Mismatch-mask Region
A natural idea to solve the mismatched region is to design an algorithm to adaptively arrange each position in the region with a proper non-inner face category (e.g., nearest neighbour classifier). However, we empirically notice that even if all mismatched regions are filled with the correct category, guided by this perfect recomposed mask, the content generated in this region tends to conflict with the original target image $T$. This is because some background information should be generated in those locations which once belonged to the face area. Such conflict would lead to disharmony at the borders of the mismatched region after Multi-Band Blending. Thus, this paper turns to train an inpainting network $P_{\tau}$ that fills the mismatched region in the blended result $\tilde{I}^{*}_{\rm{paste}}$ in the pixel-level rather than the mask-level.
The details of the face-inpainting network $P_{\tau}$ are shown in Fig. 6. We detail the training process in the Appendix. During inference, given the pasted $\tilde{I}_{\rm{paste}}^{*}$ and the driven source $I_{\rm{d}}$, the mismatch regions $M_{\rm{r}}$ of these two faces is calculated by Eqn. 17. Then the inpainting network $P_{\tau}$ recovers the erased content, producing a refined result sharing the consistent face shape with the source. Then the area ratio (ranging in $[0,1]$) of mismatch regions (white parts of $M_{\rm{r}}$ in Fig. 6) is fed to an MLP, extracting a vector $\alpha$ which modulates the generation process of the decoder. We only copy the inpainted pixels lying in mismatched regions to the input. Consequently, the network $P_{\tau}$ is capable of preserving the facial shape of the source without additionally seeking a perfect recomposed mask $M_{\text{swap}}$.
3.4 Training Objective

### v3

@@ III-B Disentangling Shape and Texture
III-B1 Mask-guided Style Extraction
Given an image $I$ and its segmentation mask $M$, we first leverage a multi-scale encoder $F_{\phi}$ to take the image $I$ as input and produce feature maps at different levels:
@@ III-B1 Mask-guided Style Extraction
where $s_{j}$ indicates the style codes of the $j$-th facial region. Then, the generator $G_{\theta}$ takes as input the style codes and the mask $M$ are fed into the generator to produce the navie swapping face $\tilde{I}$. Formally, we define $s\in\mathbb{R}^{C\times 18\times 512}$ as the proposed $\mathcal{W}^{r+}$ space.
III-B2 Mask-guided Style Injection
The original StyleGAN contains a style-based generator taking 18 style codes with the dimension of 512 as input. The style codes are used to manipulate the feature maps of the 18 intermediate layers. As shown in Fig. 4 (a), the generator, consisting of a serial of style blocks, takes a constant feature map with the spatial size of 4$\times$4 as input. Each style block includes a modulation, a demodulation, 3$\times$3 convolution layer, and a noise layer B to increase diversity. Here, the learnable kernel weights and bias in each block are denoted as $W$ and $b$. $W$ would be scaled by its corresponding style code $s$ with the shape of 1$\times$512 before the convolution layer. An additional upsampling layer by the factor of two is adopted between every two style blocks, increasing the feature resolution.
Figure 4: The comparison of the original StyleGAN and the proposed mask-guided StyleGAN. (a) The original StyleGAN contains consecutive convolution blocks. Each block contains a modulation, a demodulation, and a convolution layer. $W$ and $b$ denote the learnable kernel weights for each block, and $s$ denotes the style code. B is noise broadcast operation. An upsampling layer is used between every two blocks. (b) Our mask-guided StyleGAN regionally extends the convolution block. We sum up the intermediate feature maps of each region using its segmentation mask which is downsized in advance.
@@ III-B2 Mask-guided Style Injection
where $F_{l-1}$ and $F_{l}$ indicate the input and output feature maps of $l$-th layer. $W^{\prime}_{jl}$ represents the scaled kernel weights for the $j$-th component in the $l$-th layer, and * denotes the convolution operation. Similar to Eq. 5, the $\lfloor M\rfloor_{l}$ is the downsized mask corresponding to the $l$-th layer. Following the same modulation and demodulation as the original StyleGAN, we extend the style modulation regionally. In Eq. 8, $W_{l}$ denotes the original kernel weights for the $l$-th layer, and the $s_{jl}$ indicates the style code of $j$-th component for the $l$-th layer.
@@
Figure 5: Illustration of our re-coloring network $B_{\psi}$. During training, we first obtain the random re-colored $I^{\prime}_{\rm{AG}}$ and flipped $I^{\prime}_{\rm{A}}$ and extract their features $f_{\rm{AG}}$ and $f_{\rm{A}}$ by FPN [53]. For each region $r$, we calculate an attention map between $f_{\rm{AG}}$ and $f_{\rm{A}}$, which is then multiplied with the downsampled $I^{\prime}_{A}$ and composed to generate color guidance $\Pi$. Finally, a U-Net takes as input the concatenated $\Pi$, the input images $I^{\prime}_{\rm{AG}}$, $I^{\prime}_{\rm{A}}$ and their segmentation masks, generating the re-colored result $\tilde{I}^{*}_{\rm{rec}}$, which is used to calculate the reconstruction loss with the flipped $I^{\prime}_{\rm{A}}$. During testing, we adopt a same scheme to transfer the color from the target $I_{\rm{t}}$ to the naive swapped face $\tilde{I}$, resulting in the re-colored swapped result $\tilde{I}^{*}_{\rm{rec}}$.
III-C Lighting Transfer and Shape Inpainting
@@
Although our E4S framework provides a unique face swapping insight (i.e, disentangling shape and texture), there are still some situations that it cannot handle well. (i) When the lighting difference between source and target is huge, reconstructing the face of source in the ambient lighting of target would produce disharmonious results. (ii) During recomposing the mask $M_{\text{swap}}$, there would inevitably be some unmatched mask regions. For instance, for those positions belong to the inner face of the target, but not to that of the source, to maintain the shape of the source, these unmatched face regions should be injected with the style code of non-inner face category (hair, neck, etc.). However, since different unmatched regions may correspond to different injected style codes, it is difficult to arrange them in advance. Furthermore, even if the mismatch regions are injected with the correct style code and yield a perfect recomposed mask $M_{\text{swap}}$, guided by this mask, the content generated in the mismatch regions tends to conflict with the original target image $T$. It brings difficulty on blending the naive swapped face $\tilde{I}$ and target $T$.
@@
To address the above issues, we propose two refinement techniques for more fine-grained and lighting-consistent face swapping, which consist of: transferring lighting condition and skin tones from the target to the swapped result with an additional face re-colorer $B_{\psi}$, and training an additional inpainting network $P_{\tau}$ to restore mismatch-mask regions.
III-C1 Face Re-coloring
@@
Preserving the target lighting condition is challenging, especially when the illumination of the source is leaked into the result. Disentangling the illumination requires recovering material, geometry, and lighting of a scene, making this issue ill-posed [52]. To deal with the re-lighting challenge, we reformulate the re-lighting problem as a re-coloring problem and propose to transfer the color from the target to the swapped face by our re-colorer $B_{\psi}$. Although the swapped face color should be determined by both the source skin tone and target environment lighting, we simplify the re-lighting problem by transferring the target color to the swapped face. In practice, we found this simplification leads to reasonable and admirable results, improving the result fidelity, as shown in Fig. 8.
@@
However, it is impossible to collect annotated paired data for the training of transferring the color of a reference face to the objective face. It may take great expense and huge labor to manually paint reasonable output. Thus, we turn to perform self-supervised training, where we set the face to be colored and the reference face as the same identity during training. Specifically, for a given RGB face $I_{\rm{A}}$, we convert it into a grayscale image $I_{\rm{AG}}$. As seen in Fig. 5, we first conduct random color augmentation ${\rm{CA}}$ on the grayscale image $I_{\rm{AG}}$ and a random horizontal flip augmentation ${\rm{FA}}$ on the original image $I_{\rm{A}}$ respectively:
@@
$f_{\rm{AG}}={\rm{FPN}}({I_{\rm{AG}}^{\prime}})\in\mathbb{R}^{N\times C},f_{\rm{A}}={\rm{FPN}}(I_{\rm{A}}^{\prime})\in\mathbb{R}^{N\times C},$ (10)
@@
Furthermore, since the high computational cost of calculating cosine similarity in 1024$\times$1024 resolution (e.g. calculating cosine similarity matrix between the skin parts of two 1024$\times$1024 faces requires 64GB memory in average), we train the U-Net in 256$\times$256 resolution. Although the memory is more efficient, the produced $256$ re-coloring result $\tilde{I}^{*}_{\rm{rec}}$ looks very blur when we directly resize it to 1024$\times$1024 resolution. Therefore, we use a face super-resolution method [54] to enhance the low-resolution $\tilde{I}^{*}_{\rm{rec}}$ and paste the enhanced recolored image to $\tilde{I}$ based on a low-pass mask $M_{\rm{rec}}$ where the high-frequency pixels is removed by a Sobel filter, where the final enhanced and pasted recolored output is denoted as $\tilde{I}_{\rm{rec}}$, as shown in Fig. 8.
Figure 6: Illustration of inpainting network $P_{\tau}$, which adaptively inpaints the mismatch regions on the pixel levels. Given a driven face $I_{\rm{d}}$ and the pasted $\tilde{I}_{\rm{rec}}$, we first calculate the mismatch regions mask $M_{\rm{r}}$. Then the image $\tilde{I}_{\rm{rec}}$ along with the mismatch mask $M_{\rm{r}}$ are fed to an auto-encoder to inpaint the mismatch pixels. The generation process of the decoder is modulated by the scale factor $\alpha_{1}$ and $\alpha_{2}$ extracted from the area ratio $s$ of mismatch regions. The final inpainting result is denoted as $\tilde{I}_{\rm{inp}}$. For simplicity, the normalization layers are omitted in this figure.
III-C2 Inpainting For Mismatch-mask Region
There would inevitably be quite a few mismatched regions during facial mask exchange between the driven face $M_{\rm{d}}$ and target $M_{\rm{t}}$ (Sec. III-A). Here, we divide facial mask into two categories: inner face region (eyebrows, eyes, nose, mouth, lips, face, and skin) and non-inner face region (neck, ears, hair, earrings, and background). Then, there are two types of face mismatched pixels in the recomposed mask $M_{\text{swap}}$: 1) the positions belong to the inner face of the source, but not to that of the target; 2) the positions belong to the inner face of the target, but not to that of the source. In the first case, to keep face shape of the source, we directly use face mask of the source to fill those mismatched positions. For the second case, we need to fill the position with a proper non-inner face category. In this paper, we define a mismatched region $M_{\rm{r}}$ as the second case since it cannot be solved directly:
@@ III-C2 Inpainting For Mismatch-mask Region
A natural idea to solve the mismatched region is to design an algorithm to adaptively arrange each position in the region with a proper non-inner face category (e.g., nearest neighbour classifier). However, we empirically notice that even if all mismatched regions are filled with the correct category, guided by this perfect recomposed mask, the content generated in this region tends to conflict with the original target image $T$. This is because some background information should be generated in those locations which once belonged to the face area. Such conflict would lead to disharmony at the borders of the mismatched region after Multi-Band Blending. Thus, this paper turns to train an inpainting network $P_{\tau}$ that fills the mismatched region in the blended result $\tilde{I}_{\rm{rec}}$ in the pixel-level rather than the mask-level.
The framework of our face-inpainting network $P_{\tau}$ is shown in Fig. 6. We train $P_{\tau}$ in an unsupervised manner, where we randomly erase the pixels around the face contour and then encourage the network to predict the erased regions. Specifically, given a face image $I$, we generate random mismatch masks and edit its skin belonging to these mismatch regions based on the editing ability of our RGI method, obtaining an edited face $I_{\rm{edit}}$. Then, the inpainting network $P_{\tau}$ is trained by using $I_{\rm{edit}}$ as input and $I$ as the ground-truth supervision. During inference, given the recolored $\tilde{I}_{\rm{rec}}$ and the driven source $I_{\rm{d}}$, the mismatch regions $M_{\rm{r}}$ of these two faces is calculated by Eqn. 14. Then the inpainting network $P_{\tau}$ recovers the erased content, producing a refined result sharing the consistent face shape with the source. Then the area ratio (ranging in $[0,1]$) of mismatch regions (white parts of $M_{\rm{r}}$ in Fig. 6) is fed to an MLP, extracting the scale factor $\alpha_{1}$ and $\alpha_{2}$ which guide the generation process of the decoder. Specifically, we use $\alpha_{1}$ and $\alpha_{2}$ to control the linear combination of the features extracted by the encoder and the corresponding modulated ones:
$x_{\rm{out}}=\alpha_{1}{\rm{Mod}}(x_{\rm{in}},x_{\rm{skip}})+\alpha_{2}x_{\rm{skip}},$ (15)
where the modulation process $\rm{Mod}(\cdot)$ follows the modulating convolution of StyleGAN [17], $x_{\rm{in}}$ and $x_{\rm{out}}$ denote the input and output feature of a decode layer, $x_{\rm{skip}}$ indicates the skip connection features from the corresponding layer, all the subscripts of layer index are omitted for convenience sake. We only copy the inpainted pixels lying in mismatched regions to the input. Consequently, the network $P_{\tau}$ is capable of generating the inpainting results $\tilde{I}_{\rm{inp}}$ which preserves the facial shape of the source without additionally seeking a perfect recomposed mask $M_{\text{swap}}$.
III-D Training Objective